given that tan of theta= 11/6, find the other five trigonometric functions of theta.
Answer by drk(1908)

SInce TAN (THETA) is positive, it could be in first or third quadrant. We have actually two possible sets of answers. I will assume you want only first quadrant.
FIRST QUADRANT answers:
tan is opposite over adjacent. We need hypotenuse using Pythagorean theorem as
a^2 + b^2 + c^2
or
6^2 + 11^2 = c^2
36 + 121 = c^2
157 = c^2
sqrt(157) ~ 12.53 ~ c.
sin = 11/sqrt(157)
cos = 6/sqrt(157)
tan = 11/6
csc = sqrt(157)/11
sec = sqrt(157)/6
cot = 6/11